OpenStudy (anonymous):

X Y 1 1 2 1+2 3 1+2+4 4 1+2+4+8

OpenStudy (anonymous):

Write an equation that relates X and Y

OpenStudy (anonymous):

\[y=\sum _{n=1}^x 2^{n-1} \]\[\text{Table}\left[\left\{y=\sum _{n=1}^x 2^{n-1}\right\},\{x,1,5\}\right] \]{ 1, 3, 7, 15, 31 }

OpenStudy (anonymous):

The last two lines above were from Mathematica v9 Home Edition.